OpenStudy (anonymous):

what is the domain of f(x)= (Ln x +SQRT [4-x] ) /sinx

terenzreignz (terenzreignz):

Intersect the domains of these component functions :) Because all of them have to be satisfied. What's the domain of ln(x) ?

OpenStudy (anonymous):

0 ----- 2.

OpenStudy (anonymous):

is the answeR: 0>x<4

terenzreignz (terenzreignz):

Nope :) Now, what's the domain of ln(x) ?

OpenStudy (anonymous):

idk... isnt the smallest 0? and biggest infinity?

terenzreignz (terenzreignz):

smallest is not 0, because 0 is not in the domain. more like (0 , infinity) Okay, great. Now what about the domain of \(\large \sqrt{4-x}\)

OpenStudy (anonymous):

-infinity to 4

OpenStudy (anonymous):

so isnt it just 0 to 4?

OpenStudy (jdoe0001):

\((-\infty, 4]\) I'd think

terenzreignz (terenzreignz):

Oay, yes :) Now, intersect this with the domain of ln(x) which is (0 , infinity)

terenzreignz (terenzreignz):

*okay

OpenStudy (anonymous):

how could it be -infinity .. the ln curve doesnt go that far

terenzreignz (terenzreignz):

Yes it does :D and it's +infinity and not -infinity.... negative numbers do not go inside the ln :)

OpenStudy (anonymous):

uugghhh so @terenzreignz is hte answer 0,4???!!!

terenzreignz (terenzreignz):

Important to include the () or the [] to indicate openness or closedness :)

OpenStudy (anonymous):

my -infinity q was for that other person who posted abt -inifinty to 4 being the domain

OpenStudy (jdoe0001):

hehe, ohh right, log can't take negs..

OpenStudy (anonymous):

Domain of \(\sqrt{4-x}\) = \((-\infty,4]\). Domain of \(\ln x\) = \((0,\infty)\). Find the intersection of these two domains: \[(-\infty,4]\cap(0,\infty)\]

terenzreignz (terenzreignz):

What is \[\Large (0,\infty) \cap (-\infty , 4] =\color{red}?\]

OpenStudy (anonymous):

FOR CRYING OUT LOUD IS IT (0,4)

terenzreignz (terenzreignz):

oh... I'm late lol :D

OpenStudy (jdoe0001):

hehe

terenzreignz (terenzreignz):

@liliy simply put, no :) Besides, you still have a sin(x) at the denominator :D

OpenStudy (anonymous):

dang

OpenStudy (anonymous):

okay... so:(

terenzreignz (terenzreignz):

The intersection should be \[\large (0,4]\] because the 4 is included.

OpenStudy (anonymous):

sin(x) =0 when x=2pi

OpenStudy (jdoe0001):

so, there are really 3 restrictions on the domain, 2 on the numerator and 1 on the denominator

terenzreignz (terenzreignz):

Yup, what @jdoe0001 said :D and since our tentative domain is already (0,4] then we really couldn't care less about 0 or 2pi, since... \[\Large 2\pi \approx 6.28\] which is already well beyond the boundaries of (0,4] and 0 simply is not included in (0,4] XD

OpenStudy (jdoe0001):

\((-\infty, 4]\) restriction doesn't apply however to the ln() function, so the infinity holds

terenzreignz (terenzreignz):

However, there IS nother value which would make sin(x) equal to zero, that is well within (0,4], have you spotted it yet, @liliy ? :)

terenzreignz (terenzreignz):

another*

OpenStudy (anonymous):

pi

OpenStudy (jdoe0001):